using mathematical induction prove that (n+1)^2<2n^2 for all natural numbers n=3 or greater than 3
1
Expert's answer
2020-09-30T20:08:47-0400

Step 1. Prove the result is true for "n=3":

Left Side: "(n+1)^2" Right Side: "2\u22c5n^2"

Left Side: "(3+1)^2" Right Side: "2\u22c53^2"

Left Side: "4^2" Right Side: "2\u22c59"

Left Side: "16" Right Side: "18"

Left Side "<" Right Side

Hence the result is true for "n=3".

Step 2. Assume the result is true for "n=k", where "k" is a positive integer "\geqslant3":

"(k+1)^2<2k^2"

Step 3. Prove the result is true for "n=k+1":

"(k+1+1)^2<2\u22c5(k+1)^2"

"(k+2)^2<2\u22c5(k+1)^2"

"k^2+4k+4<2\u22c5(k^2+2k+1)"

"k^2+4k+4<2k^2+4k+2"

"2<k^2" by "k\geqslant3"

So, "(k+1+1)^2<2\u22c5(k+1)^2" holds by "n=k+1", therefore

"(n+1)^2<2n^2" is true for "n\geqslant3".